OpenStudy (mendicant_bias):

I was leafing through the beginning pages of my Calculus book and just found something that kind of boggles me, a simple domain and range problem that I just don't get. (Problem below momentarily.)

OpenStudy (mendicant_bias):

The short bit of the problem that I care about is finding the domain and range of f(x), where f(x) = x. The thing that seems strange is the answer given for the domain is \[D _{f} : -\infty < x < \infty \]

OpenStudy (mendicant_bias):

But, this seems odd; why is it that, as opposed to \[D _{f}: (-\infty , \infty)\]?

OpenStudy (mendicant_bias):

Oh. Wait a sec...is this just different notation for the exact same thing?

OpenStudy (mendicant_bias):

*And even if* that is the case, why the different notation? Is this kind of like how we use the obelus and two others primarily (can't remember the name) to represent division?

ganeshie8 (ganeshie8):

yep both notations represent same thing. the second one is in interval notation, while the first one is inequality

OpenStudy (mendicant_bias):

That's really odd. I apparently totally forgot that inequalities were used to represent ranges. Thanks.

ganeshie8 (ganeshie8):

good luck with ur revision wid calculus !